In Physics, the mean was 70 and the standard deviation was 8. In Chemistry, the mean was 68 and the standard deviation was 10. If Alison scored 75 in both PHysics and Chemistry, in which test did she perform better? Justify your answer. Which mark obtained in Chemistry is equivalent to a mark of 62 in PHysics? What percentage of students scored a mark above 88 in Chemistry?
Thanks!!!

Answer by rothauserc(4718)

Allison scored 75 on both tests
:
Physics (75 - 70) / 8 = 0.625
Chemistry (75 - 68) / 10 = 0.700
:
***********************************************************
Allison performed better on the Chemistry test, since
her score is 0.7 of one standard deviation which is higher
than her performance on the Physics test
***********************************************************
:
62 in physics is (62 - 70) / 8 = -1
:
58 in chemistry is (58 - 68) / 10 = -1
:
**********************************************
58 in Chemistry is equivalent to 62 in Physics
**********************************************
:
Probability(P) ( X > 88 ) = 1 - P ( X < 88 )
:
z-score = ( 88 - 68 ) / 10 = 2
:
consult table of z-values for associated probability
:
P ( X < 88 ) = 0.9772
:
**********************************
P ( X > 88 ) = 1 - 0.9772 = 0.0228
**********************************
: